Time Frequency Brown Bag Seminar

Wednesday, April 22, 1998

12:00 noon

EQuad E415

Speaker: Anja Feldmann, Anna Gilbert and Walter Willinger

Title: Scaling analysis of random cascades with applications to network traffic

Abstract:

Recent studies have demonstrated that measured data network traffic exhibits locally complex irregularities, consistent with multifractal behavior. The presence of well-defined protocol hierarchies in modern data networks suggests that cascades of multiplicative processes provide a plausible physical explanation of the observed multifractal structure of measured network traffic. We investigate a class of multiplicative processes that serves as a cascade paradigm for and is motivated by the networking application. We present a wavelet-based time/scale analysis of these cascades to determine rigorously their global and local scaling behavior. In particular, we prove that for the class of multifractals generated by these cascades the multifractal formalism applies and is valid. Finally, we illustrate how to use these techniques in practice by analyzing a set of wide-area traffic traces, and discuss some open problems.